OpenStudy (anonymous):

how do i expand the equation into an infinite series: -(1/√(1-x^2)) dx the answer should be -(1-x^2/2+3x^4/8+5x^6/16+...) if you can show how we get the first three terms i would really appreciate thanks.

OpenStudy (anonymous):

the function was d theta/dx and dx was taken to the other side after being implicitly differentiated.

OpenStudy (anonymous):

the function after differentiated implicitly dtheta/dx=-1/sqrt 1-x^2 then dtheta=1/sqrt1-x^2

OpenStudy (anonymous):

sorry dx should be at the end

OpenStudy (loser66):

@.Sam.

OpenStudy (anonymous):

hi am only receiving this message @.Sam. and when i open the link its leading to another question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is the question which has been answered: hey I need some chain rule help *advanced chain rule* f(x)=x^5sec(1/x) answer is: -x^3sec(1/x)tan(1/x)+5x^4sec(1/x) I need to knw how to get the answer. This is so I can study for my Ap Calc AB final on monday. PLEASE HELP!! Thanks.

OpenStudy (anonymous):

and the question has been answered

OpenStudy (experimentx):

\[ \frac{1}{\sqrt{1 - 4x^2}} = \sum_{n=0}^\infty \binom{2n}{n} x^n\]

OpenStudy (experimentx):

make appropriate substitution

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how would we make the first substitution

OpenStudy (experimentx):

x^2 = 4 (x/2)^2 1/sqrt(1-x^2) = summation of 2nCn (x/2)^n

OpenStudy (anonymous):

ok am assuming the n is the first term

OpenStudy (experimentx):

\[ \frac{1}{\sqrt{1-x^2}} = \sum_{n=0}^\infty \binom{2n}{n} (\frac{x}{2})^2\] that 2nCn is called central binomail coefficient and the LHS is called generating function. http://planetmath.org/centralbinomialcoefficient

OpenStudy (experimentx):

woops!! sorry .. that's wrong. you need to change x to x^2 \[ \frac{1}{\sqrt{1-x^2}} =\frac{1}{\sqrt{1-4(x^2/4)}} = \sum_{n=0}^\infty \binom{2n}{n} (\frac{x^2}{4})^n \]

OpenStudy (anonymous):

so one final thing how would we obtain the first two terms -(1+x^2/2+3x^4/8

OpenStudy (experimentx):

put the value of n=0, 1, 2, ... so on. (2*0)!/(0!)^2 (x/4)^0 + (2*1)!/(1!)^2 (x/4)^1+(2*2)!/(2!)^2 (x/4)^3 + .... most likely, you are supposed to use binomial theorem.

OpenStudy (experimentx):

using Binomail expansion you can evaluate it as \[ (1 - x^2)^{-1/2} = 1 + (-1/2) (-x^2) + (-1/2)(-1/2-1)/2! (-x^2)^2 +\\ (-1/2)(-1/2-1)(-1/2-2)/3! (-x^2)^3 + ... \]